OpenStudy (anonymous):

(x^-4y^0z^2)^-2 a) x^8/y^2z^4 b) 1/x^8z^4 c) x^8/z^4 d) x^6y/z^4

OpenStudy (anonymous):

I solved the equation to c) x^8/z^4 Is this correct?

OpenStudy (anonymous):

oh wait.....by using exponent rule for negative exponents...If "a" is a real number other than "0" and "n" is an integer, then, a^-n=1/a^n and 1/a^-n=a^n. I attempted the equation again and solved to b) 1/x^8z^4

jhonyy9 (jhonyy9):

x^-4 = 1/x^4 y^0 =1 z^2 =z^2 so than your exercise will be ( (z^2)/(x^4) )^-2 so than 1 1 x^8 --------------- = ----------- = ------ ((z^2)/(x^4))^2 (z^4)/(x^8) z^4 hope so much that is understandably good luck bye

jhonyy9 (jhonyy9):

c. is right sure

OpenStudy (anonymous):

Thank you!

jhonyy9 (jhonyy9):

yw good luck bye